The present value of an annuity for three years is 5000, the payments are made at the end of every six months, and the interest rate is 6% compounded monthly. How large is each payment?
1
Expert's answer
2021-09-07T10:08:19-0400

"FV=5000 \\
i=0.06 \times 6 = 0.36 \; \% \; per \;one \; period \\
n=3 \times 2=6 \\
FV = C \times [\frac{1-(1+i)^{-n}}{i}] \\
C= \frac{FV}{[\frac{1-(1+i)^{-n}}{i}]} \\
= \frac{5000}{\frac{1- (1+0.36)^{-6}}{0.36}} \\
= \frac{5000}{2.338} \\
=2137.86"

Each payment is 2137.86